OpenStudy (anonymous):

find the exact trigonometry value of sin(15),sin(75),cos(-15)and tan(75)

OpenStudy (anonymous):

\(\large sin(15)=sin(45-30) \rightarrow \) use difference formula for sine \(\large sin(75)=sin(45+30) \rightarrow \) use sum formula for sine \(\large cos(-15)=cos(30-45) \rightarrow \) use difference formula for cosine \(\large tan(75)=tan(45+30) \rightarrow \) use sum formula for tangent